A rectangle has a perimeter of 68 units and 2 lengths of 12 units. Find the width.





Answer by macston(5194)

P=perimeter; L=length; W=width
P=2(L+W)
68 units=2(12 units+W) Divide each side by 2.
34 units=12 units+W Subtract 12 units from each side.
22 units=W ANSWER: The width of the rectangle is 22 units.
CHECK:
P=2(L+W)
68 units=2(12 units + 22 units)
68 units=2(24 units)
68 units=68 units